# Exploring Weak measurements within the Einstein-Dirac Cosmological framework

### v1

@@ Abstract
Our study applies the Two-State Formalism alongside weak measurements within a spatially homogeneous and isotropic cosmological framework, wherein Dirac spinors are intricately coupled to classical gravity. To elucidate this, we provide detailed formulations for computing the weak values of the energy-momentum tensors, the Z component of spin, and the characterization of pure states. Weak measurements appear to be a generalization and extension of the computation already made by Finster an Hainzl, [1], in A spatially homogeneous and isotropic Einstein-Dirac cosmology. Our analysis reveals that the acceleration of the Universe expansion can be understood as an outcome of postselection, underscoring the effectiveness of weak measurement as a discerning approach for gauging cosmic acceleration.
@@ I INTRODUCTION
Measuring involves assigning numerical values to the attributes or characteristics of a phenomenon. The impact of the observer on measurement results is a complex consideration in scientific research. In quantum mechanics, the alteration of a quantum state after its measurement is well known, but its interpretation is not obvious. For example, the famous double-slit experiment demonstrates that particles like electrons can behave as particles and waves, and measurement prescribes their behavior [2, 3].
@@
Weak measurements offer a procedure to measure the system state that minimizes its disturbance with minimal perturbation, achieved through a weak coupling between the measurement apparatus and the system’s state. Notably, these measurements enable the characterization of a particle’s trajectory in the double-slit experiment without disrupting the interference pattern, as demonstrated in previous studies [3, 4, 5, 6]. Unlike classical or ideal measurements, which result in the stochastic collapse of the system state into one of the eigenstates of the measured observable [7], weak measurements do not induce a collapse of the state vector. Instead, they introduce a small angular bias to the state vector, and the measurement device exhibits a superposition of multiple values rather than a clear eigenvalue [4]. Consequently, weak measurements unveil unconventional weak values, including complex numbers.
@@
In literature, some authors have explored weak measurement and Two-state vectors formalism within the theoretical framework of cosmology and ontology. George Ellis and Rotman [8] proposed a paradigm that challenges conventional notions of time and reality, termed the Crystallizing Block Universe (CBU), an extension of the Emergent Block Universe (EBU). According to their perspective, past, present, and future coalesce, amalgamating space-time into a singular entity. The future is conceptualized as a superposition of myriad possibilities, while the past remains immutable. The transition between these temporal states predominantly unfolds in the present, albeit in a non-uniform manner. Within the CBU framework, certain discrete patches of quantum indeterminacy endure and are resolved later on. Furthermore, CBU models integrate the Two-time interpretation [9] of quantum mechanics. Nevertheless, this theoretical and ontological paradigm lies beyond the purview of our current study.
@@
Davies [10] is the first to propose applying weak measurements theory combined with pre-and-post-selection to quantum cosmology and to explore the potential large-scale cosmological effects arising from this new sector of quantum mechanics. He illustrated the theory with a two-spacetime dimensional toy model of a scalar field with mass $m$ propagating in an expanding Universe with the scale factor $a(t)$ and metric $ds^{2}=dt^{2}-a(t)dx^{2}$. He resolved a debate regarding the coupling of pre- and postselected quantum fields to the gravity and proposed an experimental test for weak values in cosmology. He observed that in the literature, in equation $G_{\mu\nu}+\text{higher order terms in curvature }=\frac{\langle 0_{fin}|T_{\mu\nu}|0_{in}\rangle}{\langle 0_{fin}|0_{in}\rangle}$ originally derived by DeWitt, who adapted the Schwinger effective action theory of quantum electrodynamics to the gravitational case, the source term is nothing else than the weak values of the stress-energy-momentum tensor $T_{\mu\nu}$, [11]. This equation was widespread in 1970, [12], [13], [14].
@@
We aim to examine weak measurements of Dirac particles within the framework of time symmetry as applied to the Einstein-Dirac system in a homogeneous and isotropic space such as the Friedman-Lemaître-Robertson-Walker( FLRW) space. To the best of our knowledge, no prior research has explored this direction. Drawing upon insights from the paper of Finster and Hainzl, titled A Spatially Homogeneous and Isotropic Einstein-Dirac Cosmology [1], we endeavor to extend certain findings to the domain of weak measurements. Among the outcomes of our investigation are the computation of weak values, including those pertaining to energy-momentum, pure states, and the Z component of spin. Additionally, we demonstrate that the epochs of accelerated expansion in the Universe may be regarded as consequences of postselection. This corroborates earlier studies that have explored the potential of spinor fields to elucidate phenomena such as the inflationary period in the early universe and, subsequently, the concept of dark energy. Notably, Anastopoulos arrived at a similar conclusion, albeit without invoking a spinor field or dark energy, thus emphasizing alternative avenues for understanding the observed cosmological acceleration. Ribas reached the same conclusion without resorting to weak measurements or postselection techniques. References to relevant literature supporting these assertions include [16, 17, 18, 19, 20, 21].
In the time-symmetric formulation of quantum mechanics [22], a quantum system is not characterized by a single state vector but by Two-State vectors at a specific time. The preselected state is determined by measurements conducted on the system at times prior to $t_{0}$, while the postselected state by measurements at times $t>t_{0}$. The postselected state is conceptualized as a quantum state evolving backward in time. Weak measurements occur in the interval between these conventional measurements through a weak coupling between the measurement apparatus and the system [4].
@@
$\hat{P}|\psi_{w}\rangle$ $=|\psi_{fin}\rangle\langle\psi_{fin}|e^{(-i\hat{H}T)}|\psi_{in}\rangle\otimes|\phi(x)\rangle$
Assuming that $\hat{P}_{d}$ is distributed around 0 with low variance, its conjugate variable $\hat{Q}_{d}$ will have high variance, and the measurement will be weak. Hence, we shall get:
$|\psi_{fin}\rangle\langle\psi_{fin}|(1-i\hat{A}\otimes\hat{P}_{d}\cdot T)|\psi_{in}\rangle\otimes|\phi(x)\rangle$
$=|\psi_{fin}\rangle\otimes\langle\psi_{fin}|\psi_{in}\rangle e^{-i\left(\langle\hat{A}\rangle_{w}\hat{P}_{d}T\right)}$
where $\langle\hat{A}\rangle_{w}=\frac{\langle\psi_{fin}|\hat{A}|\psi_{in}\rangle}{\langle\psi_{fin}|\psi_{in}\rangle}$ is the so-called weak measurement of the observable $\hat{A}$. The probabilities to get this result is $|\langle\psi_{fin}|\psi_{in}\rangle|^{2}$.
Weak values are weak measurement results and must be understood as statistical averages. They are also members of a decomposition of eigenvalues weighted by the relative probabilities of finding the out states among the ensemble of states vectors identically prepared in $|\psi_{in}\rangle$. Weak measurement should then be understood as a conditional probability where one of the events undergoes a perturbation. Dressel [24] clarifies this probabilistic understanding of weak values. He shows that weak values characterize the relative correction to a detection probability $|\langle\psi_{fin}|\psi_{in}\rangle|^{2}$ due to a small intermediate perturbation $\hat{U}(\epsilon)$ that results in a modified detection probability $|\langle\psi_{fin}|\hat{U}(\epsilon)\psi_{in}\rangle|^{2}$. Alternatively, in a simple way, he defines weak values as complex numbers that one can assign to the powers of a quantum observable operator $\hat{A}$ using two states: an initial state $|\psi_{in}\rangle$ and a final state $|\psi_{fin}\rangle$. Weak values might lie outside the spectrum of eigenvalues and can be complex numbers. One interesting phenomenon is the amplification obtained when the preselected state and the postselected one are almost orthogonal [4, 10].
@@
For the Einstein-Dirac in FLRW context, the preselected state will be the spinor solution of the Einstein-Dirac equation in the limit of the massless Universe, that is, the radiation Universe, whereas the final state will be the spinor solution of the Einstein-Dirac equation for the dust Universe. The implementation of weak measurement in this specific context is intended to acquire geometrical information about the Universe. The subsequent computational analysis is made feasible by the association between weak measurement and the Berry phase [7]. This has never been explored before, as far as we know.
@@
Weak measurements present numerous advantages, foremost among them being the capability to detect exceedingly subtle effects while minimizing disturbance to the system state. Onur Hosten and Paul Kwiat [26] exemplified the utility of weak measurement techniques in amplifying the spin Hall effect, Meng-Jun Hu [27] discussed in his paper the proposal for Weak Measurements Amplification based Laser Interferometer Gravitational-wave Observatory (WMA-LIGO) to detect gravitational waves using weak measurements to amplify ultra-small phase signals, while Dixon et al. [28] applied these methods to enhance the detection of minute transverse deflections in an optical beam. This facet of weak measurements proves highly advantageous in cosmology, enabling the amplification and measurement of wave functions originating in the remote past, close to singularities.
This paper is organized as follows: After the introduction, we give a summary of the Einstein-Dirac system and then find the spinor solutions for the radiation Universe and the dust one. In section III, we give the results of our main computations on weak measurements on the Einstein-Dirac solutions.In section IV, we give a conclusion. The appendices provide some details of our computations.
@@ II The Einstein-Dirac Solution in Friedmann-Lemaître-Robertson-Walker Space
where $\kappa$ represents the gravitational constant, $T^{i}_{j}$ the energy-momentum tensor of the Dirac particles, $\mathcal{D}$ the Dirac operator, $m$ the mass of the Dirac particles, and $\Psi$ the Dirac wave function.
@@
where $\sigma^{r},\sigma^{\theta}\text{and}\sigma^{\phi}$ are linear combinations of the Pauli matrixes defined as follows
@@
and the function
$\begin{array}[]{lll}f(r)&=&\sqrt{1-r^{2}},\hskip 28.45274ptr\in(0,1)\\ \end{array}$ (11)
@@
Here, $\xi$ represents the spinor $\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$, $\sigma^{i}$ are the Pauli matrices, and $\vec{e}_{i}$ are the standard basis vectors in $\mathbb{R}^{3}$ ($i\in\{1,2,3\}$).
@@
Within weak measurement and the Two Vectors State Formalism this research primarily explores solutions to the Einstein-Dirac equation under specific conditions. These conditions involve both the mass parameter ($m$) and the coupling constant ($\lambda$) approaching zero. In this regime, we distinguish between preselected states and postselected states, the solutions of (34) for respectively $m=0$ and $\lambda=0$.
In the case where $m=0$, the Einstein-Dirac equation simplifies to the well-known FLRW equation, describing the Universe’s behaviour in the radiation-dominated phase. Conversely, in the second scenario, where $\lambda=0$, corresponds to a dust-dominated Universe. Notably, for sufficiently large values of the scale factor $R$, the Universe exhibits classical behaviour similar to the dust-dominated case.
The above assumption could be seen easily from the complete and classical Friedmann general equation with $k=1,0,-1$ corresponding to closed, flat and opened universe respectively
@@
where $\rho_{m}$ is the density of matter, $\rho_{r}$ is the density of radiation and $\Lambda$ is the vacuum energy. Just after the Big Bang, the Universe was too small. For $R\ll 1$, whatever the terms in the right side of the expression (37), the radiative one would dominate in the evolution of the Universe. The expression becomes
@@
While, ignoring the vacuum energy, if $R\gg 1$, the mass term would dominate and the expression (37) becomes
@@
Figure 1: Preselection and Postselection à $t_{in}$ and $t_{out}$
@@
We set the following conditions. The expression 44 must very that at a certain time $t=t_{B},\frac{dR}{dt}|_{t_{B}}=0\,\,\text{and}R(t_{B})=R_{0}$. This implies that $w^{1}=-\frac{R^{2}_{0}}{\lambda}$. The meaning is that, at time $t_{B}$, the radiative energy ceases to be the driving force and the matter takes over. Of course, the solution is a simplification of the big picture. The solutions to the differential scale function equation is
@@
$\vec{w}(0)=\left(-\frac{R^{2}_{0}}{\lambda},\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\sin(\phi_{0}),\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos(\phi_{0})\right),$ (46)
@@
$\left\{\begin{array}[]{lll}{w}^{1}&=&-\frac{R^{2}_{0}}{\lambda},\\ {w}^{2}&=&-\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\sin(4\lambda G(t)),\\ {w}^{3}&=&\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos\left(4\lambda G(t)\right),\end{array}\right.$ (50)
where $G(t)=\tan^{-1}\left(\frac{t}{R(t)}\right)$ and $\phi_{0}=0$ We can notice that the Bloch vector is a function of the scale function $R(t)$.
@@
Replacing $\alpha$ by $\rho_{in_{1}}e^{i\eta_{in_{1}}}$ and $\beta$ by $\rho_{in_{2}}e^{i\eta_{in_{2}}}$ into (54), where $\rho$ is the modulus of $\alpha$ and $\eta$ its argument. We differentiate modulus and arguments from preselected and postselected states with indices. We obtain
@@
$\eta_{in_{1}}-\eta_{in_{2}}$ $=$ $-\arctan\left(\frac{\sqrt{\lambda^{2}N^{2}-R^{4}_{0}}\sin(4\lambda G(t))}{R^{2}_{0}}\right).$ (59)
Also from (58) and using the norm from equation (36), we obtain
$\rho_{in_{1}}$ $=\frac{\sqrt{2}}{2}\sqrt{N+\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos\left(4\lambda G(t)\right)},$ (60)
$\rho_{in_{2}}$ $=\frac{\sqrt{2}}{2}\sqrt{N-\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos\left(4\lambda G(t)\right)}.$ (61)
@@
For the scale function to be real, $R(t_{in})$ must be real. This implies that $0\leq t_{\text{in}}\leq R_{0}$. Within this period, the scale function increases with time. In our approximation the scale function stops to increase at $R_{0}$. But in the general picture, the driving force due to matter takes over and the Universe continues to expand. Let us point out also that the period within the radiation-dominated Universe is very short compared to the lifespan of the Universe. This period is approximated as $72000$ years vs $13,9$Gyrs today [31]. We also remind that within this period, the scale function is very small. The radiative period ceases where the scale function R approximately attains the value $R\approx 3.10^{-4}$. Finster in [1] shows a bouncing behaviour of the scale function around singularity.
We can illustrate the behaviour of the scale acceleration in the presence of the fermions field. This graphic is constructed only for a Universe without matter.
Figure 4: The scale acceleration in the Universe dominated by fermions.
@@
$\left\{\begin{array}[]{lll}\dot{w}^{1}&=&0\\ \dot{w}^{2}&=&-2m{w}^{3}\\ \dot{w}^{3}&=&2m{w}^{2},\end{array}\right.$ (65)
@@
We differ from Finster in [1] and we set that at $t_{max},\frac{dR}{dt}|_{t_{max}}=0$, and $R(t_{max})=R_{max}$. This implies that $w^{1}=-\frac{R_{max}}{m}$. One easily sees that the two differential equations decouple. The block vector solution will not depend on the scale function. These equations yield the following solutions
$R(g)$ $=$ $\frac{R_{max}}{2}\left(1-\cos(g)\right),$ (67)
$t(g)$ $=$ $\frac{R_{max}}{2}\left(g-sin\left(g\right)\right),$ (68)
where g is a parameter varying between 00 and $2\pi$.
Figure 5: This parametric solution is set for $R_{max}=10$. This value is taken from Finster in [1]. The time is derived from the formula $t(g)=\frac{R_{max}}{2}\left(g-sin\left(g\right)\right)$.

### v2

@@ Abstract
Our study applies the Two-State Formalism alongside weak measurements within a spatially homogeneous and isotropic cosmological framework, wherein Dirac spinors are intricately coupled to classical gravity. We compute the weak values of the energy-momentum tensors, the Z component of spin, and the pure states. Weak measurements are a generalization and extension of the computation already made by Finster and Hainzl [1] in A spatially homogeneous and isotropic Einstein-Dirac cosmology. Our analysis reveals that the acceleration of the Universe expansion can be understood as an outcome of post-selection, underscoring the effectiveness of weak measurement as a discerning approach for gauging cosmic acceleration.
@@ I INTRODUCTION
Measuring involves assigning numerical values to the attributes or characteristics of a phenomenon. The impact of the observer on measurement results is a complex consideration in scientific research. In quantum mechanics, the alteration of a quantum state after its measurement is well known, but its interpretation is not obvious. For example, the famous double-slit experiment demonstrates that entities like electrons can behave as particles and waves, and measurement prescribes their behaviour [2, 3].
@@
Weak measurements offer a procedure for measuring the system state through a weak coupling between the measurement apparatus and the system state, leaving the measured system almost undisturbed [4]. The paid cost is that the pointer variable is not sharp but instead has a broad spread. Weak measurements have enabled, for instance, the characterization of a particle’s average trajectory in the double-slit experiment without disrupting the interference pattern [3, 5, 6, 7]. Unlike classical or ideal measurements, which result in the stochastic collapse of the system state into one of the eigenstates of the measured observable [8], weak measurements do not induce a collapse of the state vector. Instead, they introduce a small bias to the state vector, and the measurement device exhibits a superposition of multiple values rather than a clear eigenvalue [5]. Consequently, weak measurements unveil unconventional weak values, including complex numbers.
@@
In the literature, a few authors have explored weak measurements and the Two-state vector formalism within the theoretical framework of cosmology and ontology. George Ellis and Rotman [9] proposed a paradigm that challenges conventional notions of time and reality, termed the Crystallizing Block Universe (CBU), an extension of the Emergent Block Universe (EBU). According to their perspective, past, present, and future coalesce, amalgamating space-time into a singular entity. The future is conceptualized as a superposition of a myriad of possibilities, while the past remains immutable. The transition between these temporal states predominantly unfolds in the present, albeit in a non-uniform manner. Within the CBU framework, certain discrete patches of quantum indeterminacy endure and are resolved later on. Furthermore, CBU models integrate the Two-time interpretation [4] of quantum mechanics. Nevertheless, this theoretical and ontological paradigm lies beyond the purview of our current study.
@@
Davies [10] was the first to propose applying weak measurements theory combined with pre- and post-selection to quantum cosmology and to explore the potential large-scale cosmological effects arising from this new sector of quantum mechanics. He illustrated the theory with a two-dimensional spacetime toy model of a scalar field with mass $m$ propagating in an expanding Universe with the scale factor $a(t)$ and metric $ds^{2}=dt^{2}-a(t)dx^{2}$. He resolved a debate regarding the coupling of pre- and post-selected quantum fields to gravity and proposed an experimental test for weak values in cosmology. He observed that, in the literature, the equation $G_{\mu\nu}+\text{higher order terms in curvature }=\frac{\langle 0_{fin}|T_{\mu\nu}|0_{in}\rangle}{\langle 0_{fin}|0_{in}\rangle}$ originally derived by DeWitt, who adapted the Schwinger effective action theory of quantum electrodynamics to the gravitational case, the source term is nothing else than the weak value of the stress-energy-momentum tensor $T_{\mu\nu}$, [11]. This equation was widespread in 1970, [12, 13, 14].
@@
We aim to examine weak measurements of Dirac particles within the framework of time symmetry as applied to the Einstein-Dirac system in a homogeneous and isotropic space such as the Friedman-Lemaître-Robertson-Walker(FLRW) space. To our knowledge, no prior research has explored this direction. Drawing upon insights from the paper of Finster and Hainzl, titled A Spatially Homogeneous and Isotropic Einstein-Dirac Cosmology [1], we endeavor to extend specific findings to the domain of weak measurements. Among the outcomes of our investigation are the computations of weak values, including those about energy-momentum, pure states, and the Z component of spin. Additionally, we demonstrate that cosmic acceleration in the Universe may be regarded as a consequence of post-selection. This corroborates earlier studies that have explored the potential of spinor fields to elucidate phenomena such as the inflationary period in the early Universe and, subsequently, the concept of dark energy. Notably, Anastopoulos arrived at a similar conclusion without invoking a spinor field or dark energy, thus emphasizing alternative avenues for understanding the observed cosmological acceleration. Ribas reached the same conclusion without resorting to weak measurements or post-selection techniques. References to relevant literature supporting these assertions include [16, 17, 18, 19, 20, 21].
In the time-symmetric formulation of quantum mechanics [22], a quantum system is characterized by two state vectors at a specific time instead of a single one. The pre-selected state is determined by measurements conducted on the system prior to $t_{0}$, while the post-selected state is by measurements at times $t>t_{0}$. The post-selected state is conceptualized as a quantum state evolving backward in time. Weak measurements occur in the interval between these conventional measurements through a weak coupling between the measurement apparatus and the system [5].
@@
$|\psi_{w}\rangle$ $=e^{(-i\hat{H}T)}|\psi_{in}\rangle\otimes|\phi(x)\rangle$
$=\sum_{i}\alpha_{i}|a_{i}\rangle\otimes|\phi(q-g_{0}Ta_{i})\rangle.$
wherein$|\psi_{in}\rangle=\sum_{i}\alpha_{i}|a_{i}\rangle$ is expressed in vector basis of the observable $\hat{A}$. Making strong measurement on the device pointer state will yield $\langle q|\psi_{w}\rangle=\sum_{i}\alpha_{i}\phi(q-g_{0}Ta_{i})|a_{i}\rangle$.
Weak measurement with post-selection is obtained by making a projection with the post-selected state $\langle\psi_{fin}|$. In detail, let $\hat{P}=|\psi_{fin}\rangle\langle\psi_{fin}|\otimes I$ be an operator which projects onto the post-selected state $|\psi_{fin}\rangle$, and $|\psi_{w}\rangle=e^{(-i\hat{H}T)}|\psi_{in}\rangle\otimes|\phi(x)\rangle$ the entangled state of the system with the state of the measurement device [5],[23], then we get
$\hat{P}|\psi_{w}\rangle$ $=|\psi_{fin}\rangle\langle\psi_{fin}|e^{(-i\hat{H}T)}|\psi_{in}\rangle\otimes|\phi(x)\rangle.$
Hence, weak coupling between the system and the device will yield:
$\hat{P}|\psi_{w}\rangle\approx|\psi_{fin}\rangle\langle\psi_{fin}|(1-i\hat{A}\otimes\hat{P}_{d}\cdot g_{0}T)|\psi_{in}\rangle\otimes|\phi(q)\rangle$
$=|\psi_{fin}\rangle\langle\psi_{fin}|\psi_{in}\rangle e^{-i\left(\langle\hat{A}\rangle_{w}\hat{P}_{d}T\right)}\otimes|\phi(q)\rangle$
where $\langle\hat{A}\rangle_{w}=\frac{\langle\psi_{fin}|\hat{A}|\psi_{in}\rangle}{\langle\psi_{fin}|\psi_{in}\rangle}$ is the so-called weak value of the observable $\hat{A}$. The probabilities to get this result is $|\langle\psi_{fin}|\psi_{in}\rangle|^{2}$.
Weak values are weak measurement results and must be understood as statistical averages. They are also members of a decomposition of eigenvalues weighted by the relative amplitudes of finding the out states among the ensemble of states vectors identically prepared in $|\psi_{in}\rangle$. A weak value should then be understood as a conditional probability where one of the events undergoes a perturbation. Dressel [24] clarifies this probabilistic understanding of weak values. He shows that weak values characterize the relative correction to a detection probability $|\langle\psi_{fin}|\psi_{in}\rangle|^{2}$ due to a small intermediate perturbation $\hat{U}(\epsilon)$ that results in a modified detection probability $|\langle\psi_{fin}|\hat{U}(\epsilon)|\psi_{in}\rangle|^{2}$. Alternatively, in a simple way, he defines weak values as complex numbers that one can assign to the powers of a quantum observable operator $\hat{A}$ using two states: an initial state $|\psi_{in}\rangle$ and a final state $|\psi_{fin}\rangle$. Weak values might lie outside the spectrum of eigenvalues and can be complex numbers. One interesting phenomenon is the amplification obtained when the pre-selected state and the post-selected one are almost orthogonal [5, 10].
@@
This idea of weakness of all observations and measurements in cosmology is rooted in classical field. Dressel, in his paper [25] have shown that, given two spacetime hypersufaces $\sigma_{I}$ and $\sigma_{F}$, on which field states $|I\rangle$ and $\langle F|$ are define as boundary conditions, the classical background field $\phi$ has the form of a weak value of the quantum field operator $\phi=\frac{\langle F|\hat{\phi}|I\rangle}{\langle F|I\rangle}$. Any weak interaction, that is any weak measurement that does not appreciably perturb the classical background field or its boundary conditions will result in a quantum weak measurement whether or not the measurement coupling is weak for every field excitation.
@@
For the Einstein-Dirac in FLRW context, the pre-selected state will be the spinor solution of the Einstein-Dirac equation in the limit of the massless Universe, that is, the radiation Universe, whereas the final state will be the spinor solution of the Einstein-Dirac equation for the dust Universe. The implementation of a weak measurement in this specific context is intended to acquire geometrical information about the Universe. The subsequent computational analysis is made feasible by the association between weak measurement and the Berry phase [8]. This has never been explored before, as far as we know.
@@
Weak measurements present numerous advantages, foremost among them being the capability to detect exceedingly subtle effects while minimizing disturbance to the system state. Onur Hosten and Paul Kwiat [27] exemplified the utility of weak measurement techniques in amplifying the spin Hall effect of light, Meng-Jun Hu [28] discussed in his paper the proposal for Weak Measurements Amplification based Laser Interferometer Gravitational-wave Observatory (WMA-LIGO) to detect gravitational waves using weak measurements to amplify ultra-small phase signals, while Dixon et al. [29] applied these methods to enhance the detection of minute transverse deflections in an optical beam. This facet of weak measurements proves highly advantageous in cosmology, enabling the amplification and measurement of wave functions originating in the remote past, close to singularities.
This paper is organized as follows: after the introduction, we give a summary of the Einstein-Dirac system and then find the spinor solutions for the radiation Universe and the dust one. In section III, we give the results of our main computations on weak measurements on the Einstein-Dirac solutions. In section IV, we give a conclusion. The appendices provide more details of our computations.
@@ II The Einstein-Dirac Solution in Friedmann-Lemaître-Robertson-Walker Space
where $\kappa$ represents the gravitational constant, $R^{i}_{j}$ the Ricci tensor, $R$ the scalar curvature, $T^{i}_{j}$ the energy-momentum tensor of the Dirac particles, $\{i,j\}$ the indices representing the coordinates $\{t,r,\theta,\phi\}$, $\mathcal{D}$ the Dirac operator, $m$ the mass of the Dirac particles, and $\Psi$ the Dirac wave function.
@@
where $\sigma^{r},\sigma^{\theta}\text{and}\sigma^{\phi}$ are linear combinations of the Pauli matrices defined as follows
@@
wherein the Pauli matrices are
$\sigma^{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\sigma^{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\quad\sigma^{3}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ (7)
and the function $f(r)$ is
$\begin{array}[]{lll}f(r)&=&\sqrt{1-r^{2}},\hskip 28.45274ptr\in(0,1).\\ \end{array}$ (9)
@@
Here, $\xi$ represents the spinor $\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$, $\sigma^{i}$ the Pauli matrices, and $\vec{e}_{i}$ the standard basis vectors in $\mathbb{R}^{3}$ ($i\in\{1,2,3\}$).
@@
Within weak measurement and the Two Vectors State Formalism, this research primarily explores solutions to the Einstein-Dirac equation under specific conditions. These conditions involve either the mass parameter ($m=0$), in which case the solution of (32) is the preselected state, or the eigenvalue ($\lambda=0$), in which case the solution of the same equation is the post-selected state.
In the case where $m=0$, the Einstein-Dirac equation simplifies to the well-known FLRW equation, describing the Universe’s behaviour in the radiation-dominated phase. Conversely, in the second scenario, $\lambda=0$ corresponds to a dust-dominated Universe in the second scenario. Notably, for sufficiently large-scale factor $R$, the Universe exhibits classical behaviour similar to the dust-dominated case.
The above assumption can be linked easily from the complete and classical Friedmann general equation with $k=1,0,-1$ corresponding to closed, flat and opened universe respectively
@@
where $\rho_{m}$ is the density of matter, $\rho_{r}$ is the density of radiation and $\Lambda$ is the vacuum energy. For $R\ll 1$, in the right side of the expression (35), the radiative term dominates in the evolution of the Universe. The expression becomes
@@
While, ignoring the vacuum energy, if $R\gg 1$, the matter term dominates and the expression (35) becomes
@@
Figure 1: pre-selection and post-selection à $t_{in}$ and $t_{out}$
@@
We set the following conditions. The expression (27) must verify that at time $t=t_{B},\frac{dR}{dt}|_{t_{B}}=0\,\,\text{and }R(t_{B})=R_{0}$. This implies that $w^{1}=-\frac{R^{2}_{0}}{\lambda}$ a constant. The solution is a simplification of the big picture. The solution to the differential scale function equation is
@@
$\vec{w}(t_{B})=\left(-\frac{R^{2}_{0}}{\lambda},\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\sin(\phi_{0}),\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos(\phi_{0})\right),$ (44)
@@
$\left\{\begin{array}[]{lll}{w}^{1}&=&-\frac{R^{2}_{0}}{\lambda},\\ {w}^{2}&=&\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\sin{\left(\Phi_{in}(t)\right)},\\ {w}^{3}&=&\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos{\left(\Phi_{in}(t)\right)},\end{array}\right.$ (48)
where
$\Phi_{in}(t)=4\lambda G(t)-\lambda\pi-\phi_{0}$ (49)
$G(t)=\tan^{-1}\left(\frac{t}{R(t)}\right)$ and $\phi_{0}$ a phase. We can notice that the Bloch vector is a function of the scale function $R(t)$.
@@
Replacing $\alpha$ by $\rho_{in_{1}}e^{i\eta_{in_{1}}}$ and $\beta$ by $\rho_{in_{2}}e^{i\eta_{in_{2}}}$ into (53), where $\rho$ is the modulus of $\alpha$ and $\eta$ its argument. We differentiate modulus and arguments from pre-selected and post-selected states with indices. We obtain
@@
$\eta_{in_{1}}-\eta_{in_{2}}$ $=$ $-\arctan\left(\frac{\sqrt{\lambda^{2}N^{2}-R^{4}_{0}}\sin{\left(\Phi_{in}(t)\right)}}{R^{2}_{0}}\right).$ (58)
Also from (57) and using the normalisation condition from (34), we obtain
$\rho_{in_{1}}$ $=\frac{\sqrt{2}}{2}\sqrt{N+\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos\left(\Phi_{in}(t)\right)},$ (59)
$\rho_{in_{2}}$ $=\frac{\sqrt{2}}{2}\sqrt{N-\sqrt{N^{2}-\frac{R^{4}_{0}}{\lambda^{2}}}\cos\left(\Phi_{in}(t)\right)}.$ (60)
@@
[…]
@@
$\left\{\begin{array}[]{lll}\dot{w}^{1}&=&0,\\ \dot{w}^{2}&=&-2m{w}^{3},\\ \dot{w}^{3}&=&2m{w}^{2},\end{array}\right.$ (64)
@@
We differ from Finster in [1] and we set that at $t_{max},\frac{dR}{dt}|_{t_{max}}=0$, and $R(t_{max})=R_{max}$. This implies that $w^{1}=-\frac{R_{max}}{m}$. One easily sees that the two differential equations decouple. The Bloch vector solution will not depend on the scale function. These equations yield the following solutions
$R(\tau)$ $=$ $\frac{R_{max}}{2}\left(1-\cos{(\tau)}\right),$ (66)
$t(\tau)$ $=$ $\frac{R_{max}}{2}\left(\tau-\sin{\left(\tau\right)}\right),$ (67)
where $\tau$ is a parameter varying between 00 and $2\pi$. The parametric equations represent a cycloid.
Figure 4: This parametric solution is set for $R_{max}=10$. This value is taken from Finster in [1]. The time is derived from the formula $t(\tau)=\frac{R_{max}}{2}\left(\tau-\sin{(\left(\tau\right)}\right)$.